# OpenWorldSAM: Extending SAM2 for Universal Image Segmentation with Language Prompts

### v3

@@ 1 Introduction
Image segmentation has long been constrained to closed-vocabulary settings, where models can only recognize objects from a predefined taxonomy [long2015fcn, ronneberger2015unet, chen2017deeplab, chen2018deeplab, he2017maskrcnn, xie2021segformer, cheng2022mask2former, li2023mask]. However, real-world applications, e.g., Embodied AI [fan2024active, laina2025findanything], demand systems that can understand open-ended language descriptions (from single nouns like “pedestrian” to rich referring expressions such as “the man in a red shirt”) and segment novel objects unseen during training. This open-vocabulary segmentation problem poses two core challenges: (1) Semantic grounding – mapping free-form text to visual entities, and (2) Instance awareness – distinguishing multiple objects that match the same description.
Detection-centric methods [ghiasi2022openseg, liang2023ovseg] relied on two-stage pipelines, first detecting class-agnostic mask proposals then classifying them with vision-language models (VLMs), e.g., CLIP [radford2021clip] and ALIGN [jia2021align]. While effective, such approaches struggle with complex queries and specialize exclusively in semantic segmentation, lacking versatility. Recent generalist models [zou2023xdecoder, shen2024ape] explore unified architectures that jointly handle vision and language, allowing a single model to perform detection, segmentation, and grounding tasks. These generalist models demonstrate impressive flexibility, but they typically involve resource-intensive pre-training. The emergence of promptable segmentation models like the Segment Anything Model (SAM) [kirillov2023sam, ravi2024sam2] offered new possibilities – it introduced a paradigm shift by allowing users to segment arbitrary objects using simple visual prompts (e.g., points, boxes). Trained on an extensive dataset, these models exhibit remarkable generalization and interactive capabilities. However, they inherently lack semantic understanding. Subsequent attempts to combine SAM with large language models (LLMs) [lai2024lisa, rasheed2024glamm, xu2024ullava] achieved language awareness but at prohibitive computational costs, imposing overwhelming overhead.
@@
    OpenWorldSAM achieves state-of-the-art zero-shot performance across six benchmarks (Figure 2), setting a new standard for open-vocabulary segmentation (e.g., 60.4 mIoU on ADE20K [zhou2017ade20k]). OpenWorldSAM also acheives strong performance in referring expression segmentation (74.0 cIoU on RefCOCOg [nagaraja2016modeling]) with substantially fewer resources compared to recent models.
@@ 2 Related Work
Open-vocabulary segmentation. Recent advances in open-vocabulary segmentation have leveraged vision-language models (VLMs) [radford2021clip, jia2021align] to overcome the constraints of traditional closed-set segmentation models. Early approaches like LSeg [li2022lseg], RegionCLIP [zhong2022regionclip] and OWL-ViT [minderer2022owlvit] established a baseline by introducing a contrastive learning framework to align image embeddings with CLIP-based text embeddings for zero-shot detection/segmentation. Subsequent methods [ghiasi2022openseg, xu2022groupvit] scaled effectively by using weak supervision of large-scale images with captions (up to millions of regions) or text-only signals, enabling more flexible and broader semantic coverage. Two-stage approaches like MaskCLIP [dong2023maskclip] and OVSeg [liang2023ovseg] further refined this paradigm by generating mask proposals using MaskFormer [cheng2021maskformer] followed by CLIP-based classification, notably boosting accuracy through mask-adapted fine-tuning. Another line of works formulated this task as a visual grounding problem and established region-text fusion [li2022glip, liu2024grounding, yan2023universal, yao2022detclip]. More recently, unified architectures such as ODISE [xu2023odise], X-Decoder [zou2023xdecoder], SEEM [zou2023seem], OpenSeeD [zhang2023openseed], HIPIE [wang2023hipie], Semantic-SAM [li2023semanticsam] and APE [shen2024ape] have integrated multiple segmentation tasks into a single framework, showing significant progress towards general-purpose models, but they typically required resource intensive pre-training.
@@
Extending SAM for text-prompted segmentation. The Segment Anything Model (SAM) [kirillov2023sam, ravi2024sam2] achieved a breakthrough in promptable segmentation by training on 1 billion masks, enabling it to generate high-quality masks for visual prompts. A flurry of recent works have explored infusing SAM with semantic or language understanding to move beyond its original prompt types. Grounded-SAM [ren2024grounded-sam] is a pioneering effort that combines an open-vocabulary detector GroundingDINO [liu2024grounding] to generate bounding boxes from a text query, then feeds those boxes as prompts into SAM. FastSAM [zhao2023fastsam] matches CLIP embeddings with regions of interest. LLM-centric works [li2023refsam, rasheed2024glamm, lai2024lisa, xu2024ullava] attempt to map large LLMs or VLMs language embeddings into SAM or SAM-like decoder’s prompt latent space to enable referring expression segmentation. Among these, LISA [lai2024lisa] pioneered “mask-as-text-embedding” approach but was limited to single-object queries. LISA++ [yang2023lisa2] introduced instance awareness through additional instruction-tuning data, though it requires LLMs to explicitly enumerate objects—a computationally expensive process. EVF-SAM [zhang2024evfsam] recently demonstrated a lightweight alternative, integrating SAM with a multi-modal BEiT-3 encoder [wang2023beit] (0.7B parameters). While achieving state-of-the-art referring segmentation accuracy with minimal parameters, it remains constrained to single-object queries. Inspired by the success of EVF-SAM, we enhance SAM further into the domain of open-vocabulary segmentation, where the goal is to segment and label all objects (“things” and “stuff”) in the scene with open-set categories.
@@ 3 Methodology
Figure 3: (a) SAM takes a visual click and outputs 3 valid masks on the same person (the person, the backpack, and a backpack region) [kirillov2023sam]. It will not output masks for the person standing next to her. (b) Tie-breakers shift the queries to distinct regions, enabling simultaneous segmentation of all three “zebra” instances. (c) Naïve approach [zhang2024evfsam]: A single language query for “zebra” causes SAM2 to segment only the most salient instance.
@@
Multi-modal encoder. We leverage BEiT-3 [wang2023beit] to encode the input description into a semantic embedding. Given an image $I$ and a text prompt $T$ (e.g., a category name or a referring expression), we feed both modalities into BEiT‐3’s encoder to obtain joint visual–text embeddings. Concretely, tokens of $T$ and patch embeddings of a downsampled $I$ are concatenated and processed by BEiT‐3, yielding a set of feature vectors $\{\mathbf{f}_{\text{[CLS]}},\mathbf{f}_{1},\dots\}$. We take the classification token $\mathbf{f}_{\text{[CLS]}}$ as a compact summary denoted as $\mathbf{p}_{\mathrm{lang}}$ of the prompt conditioned on the image content.
@@
These perturbations serve two purposes: 1) Positional disambiguation: Each $\mathbf{t}_{i}$ nudges the query towards different spatial regions (Figure 3b), mimicking how human annotators might click different points to segment each zebra. 2) Instance diversity: The tie-breakers are optimized during training to maximize coverage of distinct instances, preventing query collapse. Conceptually these queries play the role of the “object queries” in DETR [carion2020detr]. Crucially, they impose segmentation distinction for the same language semantics, making positional tie-breaking a novel and key feature for OpenWorldSAM. In practice $K=20$ covers >>99% images in COCO [lin2014coco]; for larger scenes $K$ can be increased trivially.
Soft‐prompting via cross‐attention. The perturbed queries $\{\mathbf{q}_{i}\}$ interact with SAM2’s image features through a 3-layer Transformer [vaswani2017attention] in Figure 4c, which alternates self‑attention (queries talk to each other, promoting diversity) and cross‑attention (queries look at image features). Each language‑aware query is refined on‑the‑fly by cross‑attention with the frozen SAM2 features. SAM2’s image encoder follows a hierarchical vision Transformer (“Hiera” [ryali2023hiera, bolya2023window]) that outputs three features $\{\mathbf{F}_{256\times 256},\mathbf{F}_{128\times 128},\mathbf{F}_{64\times 64}\}$ with $256^{2}$, $128^{2}$, and $64^{2}$ spatial resolutions, respectively. We operate on the level-3 features with $64^{2}$ resolution as they optimally balance precision for retainaing boundary details and computational efficiency ($16\times$ cheaper than full‑resolution attention). They are also used by SAM2 for mask decoding by default [ravi2024sam2]. The soft prompting Transformer computes $\mathbf{q}^{\prime}_{i}=\mathrm{CrossAttn}(\mathbf{q}_{i},\ \mathbf{F}_{64\times 64}),\;i=1,\dots,K,$ whose key/value inputs are the flattened level-3 features $\mathbf{F}_{64\times 64}\in\mathbb{R}^{4096\times 256}$. This step grounds the language-aware queries in SAM2’s high-resolution visual features, resolving ambiguities (e.g., distinguishing adjacent zebras by stripe patterns).
@@
Training. All heavy visual (Hiera) and vision‑language (BEiT‑3) encoders are kept frozen to preserve their pre‑trained knowledge and avoid costly retraining. Only the MLP projector, tie‑breakers, and the soft prompting Transformer are learnable. For each training sample and prompt, we match the $K$ predicted masks to the ground‐truth masks of class $T$ via Hungarian matching [carion2020detr], then apply a focal loss, encouraging precise segmentation of all instances described by the prompt. The tie-breakers $\mathbf{t}_{i}\in\mathbb{R}^{d_{256}}$ are implemented as learnable parameters randomly initialized from a normal distribution. During training, the Hungarian matching loss naturally encourages each $\mathbf{t}_{i}$ to specialize in different spatial regions. Notably, this mechanism requires no explicit supervision about instance counts.
@@ 4 Experiments
Datasets and metrics. We train OpenWorldSAM on the COCO2017-Stuff [lin2014coco] dataset with panoptic annotations following X-Decoder [zou2023xdecoder]. The training set contains 104k images. We evaluate the model in a zero-shot setting on eight segmentation tasks across five diverse datasets: ADE20K-150/857 [zhou2017ade20k], PASCAL VOC-20 [everingham2011pascalvoc], PASCAL Context-59/459 [mottaghi2014pascal-context], ScanNet-20/40 [dai2017scannet], and SUN-RGBD-37 [song2015sun]. Evaluation metrics include panoptic quality (PQ), mean average precision (mAP), and mean intersection-over-union (mIoU), corresponding to panoptic, instance, and semantic segmentation tasks, respectively. For referring segmentation, we finetune on RefCOCOg UMD training split. Following prior works, we report the cumulative intersection over the cumulative union (cIoU) metric on the RefCOCOg UMD validation split.
Implementation. We implement our model in PyTorch. We initialize the visual model with the weights of SAM2-Hiera-Large [ravi2024sam2] and the VLM encoder with the weights of EVF-SAM BEIT-3-Large [zhang2024evfsam]. It is trained for 25 epochs on COCO-Stuff using the AdamW optimizer with a learning rate of 1e‑4, batch size 8, on a single NVIDIA A100 GPU. Image resolution is set to 1024 for SAM2 and 224 for BEiT-3. Number of postional tie-breaks is set to 20 for COCO dataset. Our implementation details can be found in Appendix A.
@@ 4.1 Open-Vocabulary Segmentation Evaluation Protocols and Challenges
Ambiguity of open vocabulary evaluation. Most prior open-vocabulary segmentation methods—including X-Decoder [zou2023xdecoder], OVSeg [liang2023ovseg], and MaskCLIP [dong2023maskclip]—adopt a Global-Matching protocol: for each predicted mask, a model matches it against the entire dataset vocabulary using precomputed text embeddings and selects the best-aligned class. However, this strategy can be problematic when applied to datasets like ADE20K, which contain hundreds of fine-grained and overlapping labels. As observed in OVSeg [liang2023ovseg], this leads to semantically reasonable predictions being marked incorrect under exact label matching: “The ground-truth category is ‘building’ while our model predicts ‘skyscraper’. ” This ambiguity stems from the inherent subjectivity of language: synonymous or closely related concepts may be indistinguishable in a visual context, yet only one is accepted by the ground truth. We observe similar issues in our own qualitative analysis. As shown in Figure 5, X-Decoder predictions on ADE20K-857 often produce valid but non-canonical labels (e.g., ‘road’ instead of ‘runway’, or ‘screen’ instead of ‘arcade machine’), resulting in unfair penalization.
@@ 4.2 Open-Vocabulary Segmentation Performance Analysis
MSeg (B) [lambert2020mseg] 70 (M) 33.7 32.6 19.1 * 73.4 43.4 * 29.6 33.4 * *
GroupViT (S) [xu2022groupvit] 44 (M) - - * * 52.3 22.4 * * * - *
LSeg+ (B) [li2022lseg] 112 (M) - - 18.0 3.8 * 46.5 7.8 * * - *
ZegFormer (B) [ding2022zegformer] 60 (M) - - * 8.1 80.7 * * * * - -
OpenSeg (B) [ghiasi2022openseg] 86 (M) - - 26.4 8.1 70.2 44.8 11.5 * * - *
OVSeg (B) [liang2023ovseg] 0.6 (M) - - 29.6 9.0 94.5 55.7 12.4 * * - *
MaskCLIP (L) [dong2023maskclip] 428 (M) 15.1 6.0 23.7 8.2 * 45.9 10.0 * * * *
OpenSeeD (L) [zhang2023openseed] 39 (M) 19.7 15.0 23.4 * * * * * * * *
X-Decoder-Seg+ (B) [zou2023xdecoder] 28 (M) 16.9 9.5 23.8 4.6 97.8 64.7 12.1 32.2 35.1 33.8 18.5
X-Decoder (L) [zou2023xdecoder] 38 (M) 21.8 13.1 29.6 9.2 97.7 64.0 16.1 43.0 49.5 39.5 29.7
APE-B (L) [shen2024ape] 42 (M) 26.4 23.5 29.0 9.2 95.8 58.3 21.0 * * * *
ESC-Net [lee2025effective] 451 (M) - - 41.8 18.1 98.3 65.6 27.0 * * - *
@@
Table 2: Oracle-Prompts evaluation of open-vocabulary segmentation models. We report the state-of-the-art (SOTA) model X-Decoder [zou2023xdecoder]’s performance under both evaluation protocols. Other methods are omitted either because: 1) they are not SOTA, or 2) they do not support oracle-prompts evaluation.
@@
X-Decoder (L) [zou2023xdecoder] Global-Matching (default) 29.6 9.2 97.7 64.0 16.1 43.0 29.7
@@
Zero-shot open-vocabulary transfer. OpenWorldSAM generalizes out-of-the-box to a broad set of segmentation tasks without any weight adaptation. As shown in Table 1, it achieves state-of-the-art performance across almost all datasets and evaluation metrics. Its performance consistently surpasses strong baselines such as X-Decoder and APE, despite using only 4.5M trainable parameters. On ADE20K-857, OpenWorldSAM achieves 33.1% mIoU, outperforming the previous best (X-Decoder) by +23.9 absolute points (9.2 → 33.1). On PASCAL Context-459, it achieves 47.5% mIoU, improving over APE’s 21.8% by +25.7 points, and on ScanNet-40, it reaches 55.6% mIoU, a +25.9 point improvement over X-Decoder’s 29.7%. On AP score we under-perform APE, which included extra detection datasets, e.g., Objects365 [shao2019objects365], in their training recipe for better localization.
@@
Oracle-Prompts evaluation. As SAM-style models are designed for interactive segmentation, oracle prompts closely reflect practical use cases—such as human-in-the-loop annotation, robotic object search, or dynamic UI feedback. To fairly compare with the state-of-the-art generalist model X-Decoder [zou2023xdecoder], we also evaluate it under oracle prompts: we restrict its vocabulary to the ground-truth classes for each image. As shown in Table 2, OpenWorldSAM continues to outperform even under these controlled conditions. Notably, on large-vocabulary datasets such as ADE20K-857 and PASCAL Context-459, OpenWorldSAM achieves 33.1% and 47.5% mIoU, surpassing X-Decoder by +4.0 and +5.2 points, respectively. This highlights our model’s superior language grounding ability in long-tailed, fine-grained category distributions. On smaller datasets like PASCAL Context-59 and PASCAL VOC-20, where most categories overlap with COCO, X-Decoder slightly outperforms our model (75.5% vs. 73.7% mIoU and 98.1% vs. 98.0%), suggesting it benefits more from class memorization in such settings. Moreover, Figure 5 illustrates that global matching often fails despite producing correct masks. Conditioning on oracle prompts significantly reduces this ambiguity, highlighting the robustness of our evaluation protocol and the effectiveness of prompt-based segmentation.

### v4

@@ 1 Introduction
Image segmentation has long been constrained to closed-vocabulary settings, where models can only recognize objects from a predefined taxonomy [39, 49, 5, 6, 20, 63, 9, 30]. However, real-world applications, e.g., Embodied AI [18, 25], demand systems that can understand open-ended language descriptions (from single nouns like “pedestrian” to rich referring expressions such as “the man in a red shirt”) and segment novel objects unseen during training. This open-vocabulary segmentation problem poses two core challenges: (1) Semantic grounding – mapping free-form text to visual entities, and (2) Instance awareness – distinguishing multiple objects that match the same description.
Detection-centric methods [19, 34] relied on two-stage pipelines, first detecting class-agnostic mask proposals then classifying them with vision-language models (VLMs), e.g., CLIP [44] and ALIGN [22]. While effective, such approaches struggle with complex queries and specialize exclusively in semantic segmentation, lacking versatility. Recent generalist models [80, 52] explore unified architectures that jointly handle vision and language, allowing a single model to perform detection, segmentation, and grounding tasks. These generalist models demonstrate impressive flexibility, but they typically involve resource-intensive pre-training. The emergence of promptable segmentation models like the Segment Anything Model (SAM) [23, 46] offered new possibilities – it introduced a paradigm shift by allowing users to segment arbitrary objects using simple visual prompts (e.g., points, boxes). Trained on an extensive dataset, these models exhibit remarkable generalization and interactive capabilities. However, they inherently lack semantic understanding. Subsequent attempts to combine SAM with large language models (LLMs) [24, 45, 66] achieved language awareness but at prohibitive computational costs, imposing overwhelming overhead.
@@
    OpenWorldSAM achieves state-of-the-art zero-shot performance across six benchmarks (Figure 2), setting a new standard for open-vocabulary segmentation (e.g., 60.4 mIoU on ADE20K [79]). OpenWorldSAM also acheives strong performance in referring expression segmentation (74.0 cIoU on RefCOCOg [42]) with substantially fewer resources compared to recent models.
@@ 2 Related Work
Open-vocabulary segmentation. Recent advances in open-vocabulary segmentation have leveraged vision-language models (VLMs) [44, 22] to overcome the constraints of traditional closed-set segmentation models. Early approaches like LSeg [28], RegionCLIP [78] and OWL-ViT [40] established a baseline by introducing a contrastive learning framework to align image embeddings with CLIP-based text embeddings for zero-shot detection/segmentation. Subsequent methods [19, 64] scaled effectively by using weak supervision of large-scale images with captions (up to millions of regions) or text-only signals, enabling more flexible and broader semantic coverage. Two-stage approaches like MaskCLIP [16] and OVSeg [34] further refined this paradigm by generating mask proposals using MaskFormer [10] followed by CLIP-based classification, notably boosting accuracy through mask-adapted fine-tuning. Another line of works formulated this task as a visual grounding problem and established region-text fusion [32, 37, 67, 69]. More recently, unified architectures such as ODISE [65], X-Decoder [80], SEEM [81], OpenSeeD [73], HIPIE [60], Semantic-SAM [29] and APE [52] have integrated multiple segmentation tasks into a single framework, showing significant progress towards general-purpose models, but they typically required resource intensive pre-training.
@@
Extending SAM for text-prompted segmentation. The Segment Anything Model (SAM) [23, 46] achieved a breakthrough in promptable segmentation by training on 1 billion masks, enabling it to generate high-quality masks for visual prompts. A flurry of recent works have explored infusing SAM with semantic or language understanding to move beyond its original prompt types. Grounded-SAM [47] is a pioneering effort that combines an open-vocabulary detector GroundingDINO [37] to generate bounding boxes from a text query, then feeds those boxes as prompts into SAM. FastSAM [76] matches CLIP embeddings with regions of interest. LLM-centric works [33, 45, 24, 66] attempt to map large LLMs or VLMs language embeddings into SAM or SAM-like decoder’s prompt latent space to enable referring expression segmentation. Among these, LISA [24] pioneered “mask-as-text-embedding” approach but was limited to single-object queries. LISA++ [68] introduced instance awareness through additional instruction-tuning data, though it requires LLMs to explicitly enumerate objects—a computationally expensive process. EVF-SAM [75] recently demonstrated a lightweight alternative, integrating SAM with a multi-modal BEiT-3 encoder [59] (0.7B parameters). While achieving state-of-the-art referring segmentation accuracy with minimal parameters, it remains constrained to single-object queries. Inspired by the success of EVF-SAM, we enhance SAM further into the domain of open-vocabulary segmentation, where the goal is to segment and label all objects (“things” and “stuff”) in the scene with open-set categories.
@@ 3 Methodology
Figure 3: (a) SAM takes a visual click and outputs 3 valid masks on the same person (the person, the backpack, and a backpack region) [23]. It will not output masks for the person standing next to her. (b) Tie-breakers shift the queries to distinct regions, enabling simultaneous segmentation of all three “zebra” instances. (c) Naïve approach [75]: A single language query for “zebra” causes SAM2 to segment only the most salient instance.
@@
Multi-modal encoder. We leverage BEiT-3 [59] to encode the input description into a semantic embedding. Given an image $I$ and a text prompt $T$ (e.g., a category name or a referring expression), we feed both modalities into BEiT‐3’s encoder to obtain joint visual–text embeddings. Concretely, tokens of $T$ and patch embeddings of a downsampled $I$ are concatenated and processed by BEiT‐3, yielding a set of feature vectors $\{\mathbf{f}_{\text{[CLS]}},\mathbf{f}_{1},\dots\}$. We take the classification token $\mathbf{f}_{\text{[CLS]}}$ as a compact summary denoted as $\mathbf{p}_{\mathrm{lang}}$ of the prompt conditioned on the image content.
@@
These perturbations serve two purposes: 1) Positional disambiguation: Each $\mathbf{t}_{i}$ nudges the query towards different spatial regions (Figure 3b), mimicking how human annotators might click different points to segment each zebra. 2) Instance diversity: The tie-breakers are optimized during training to maximize coverage of distinct instances, preventing query collapse. Conceptually these queries play the role of the “object queries” in DETR [4]. Crucially, they impose segmentation distinction for the same language semantics, making positional tie-breaking a novel and key feature for OpenWorldSAM. In practice $K=20$ covers >>99% images in COCO [35]; for larger scenes $K$ can be increased trivially.
Soft‐prompting via cross‐attention. The perturbed queries $\{\mathbf{q}_{i}\}$ interact with SAM2’s image features through a 3-layer Transformer [57] in Figure 4c, which alternates self‑attention (queries talk to each other, promoting diversity) and cross‑attention (queries look at image features). Each language‑aware query is refined on‑the‑fly by cross‑attention with the frozen SAM2 features. SAM2’s image encoder follows a hierarchical vision Transformer (“Hiera” [50, 3]) that outputs three features $\{\mathbf{F}_{256\times 256},\mathbf{F}_{128\times 128},\mathbf{F}_{64\times 64}\}$ with $256^{2}$, $128^{2}$, and $64^{2}$ spatial resolutions, respectively. We operate on the level-3 features with $64^{2}$ resolution as they optimally balance precision for retainaing boundary details and computational efficiency ($16\times$ cheaper than full‑resolution attention). They are also used by SAM2 for mask decoding by default [46]. The soft prompting Transformer computes $\mathbf{q}^{\prime}_{i}=\mathrm{CrossAttn}(\mathbf{q}_{i},\ \mathbf{F}_{64\times 64}),\;i=1,\dots,K,$ whose key/value inputs are the flattened level-3 features $\mathbf{F}_{64\times 64}\in\mathbb{R}^{4096\times 256}$. This step grounds the language-aware queries in SAM2’s high-resolution visual features, resolving ambiguities (e.g., distinguishing adjacent zebras by stripe patterns).
@@
Training. All heavy visual (Hiera) and vision‑language (BEiT‑3) encoders are kept frozen to preserve their pre‑trained knowledge and avoid costly retraining. Only the MLP projector, tie‑breakers, and the soft prompting Transformer are learnable. For each training sample and prompt, we match the $K$ predicted masks to the ground‐truth masks of class $T$ via Hungarian matching [4], then apply a focal loss, encouraging precise segmentation of all instances described by the prompt. The tie-breakers $\mathbf{t}_{i}\in\mathbb{R}^{d_{256}}$ are implemented as learnable parameters randomly initialized from a normal distribution. During training, the Hungarian matching loss naturally encourages each $\mathbf{t}_{i}$ to specialize in different spatial regions. Notably, this mechanism requires no explicit supervision about instance counts.
@@ 4 Experiments
Datasets and metrics. We train OpenWorldSAM on the COCO2017-Stuff [35] dataset with panoptic annotations following X-Decoder [80]. The training set contains 104k images. We evaluate the model in a zero-shot setting on eight segmentation tasks across five diverse datasets: ADE20K-150/857 [79], PASCAL VOC-20 [17], PASCAL Context-59/459 [41], ScanNet-20/40 [13], and SUN-RGBD-37 [53]. Evaluation metrics include panoptic quality (PQ), mean average precision (mAP), and mean intersection-over-union (mIoU), corresponding to panoptic, instance, and semantic segmentation tasks, respectively. For referring segmentation, we finetune on RefCOCOg UMD training split. Following prior works, we report the cumulative intersection over the cumulative union (cIoU) metric on the RefCOCOg UMD validation split.
Implementation. We implement our model in PyTorch. We initialize the visual model with the weights of SAM2-Hiera-Large [46] and the VLM encoder with the weights of EVF-SAM BEIT-3-Large [75]. It is trained for 25 epochs on COCO-Stuff using the AdamW optimizer with a learning rate of 1e‑4, batch size 8, on a single NVIDIA A100 GPU. Image resolution is set to 1024 for SAM2 and 224 for BEiT-3. Number of postional tie-breaks is set to 20 for COCO dataset. Our implementation details can be found in Appendix A.
@@ 4.1 Open-Vocabulary Segmentation Evaluation Protocols and Challenges
Ambiguity of open vocabulary evaluation. Most prior open-vocabulary segmentation methods—including X-Decoder [80], OVSeg [34], and MaskCLIP [16]—adopt a Global-Matching protocol: for each predicted mask, a model matches it against the entire dataset vocabulary using precomputed text embeddings and selects the best-aligned class. However, this strategy can be problematic when applied to datasets like ADE20K, which contain hundreds of fine-grained and overlapping labels. As observed in OVSeg [34], this leads to semantically reasonable predictions being marked incorrect under exact label matching: “The ground-truth category is ‘building’ while our model predicts ‘skyscraper’. ” This ambiguity stems from the inherent subjectivity of language: synonymous or closely related concepts may be indistinguishable in a visual context, yet only one is accepted by the ground truth. We observe similar issues in our own qualitative analysis. As shown in Figure 5, X-Decoder predictions on ADE20K-857 often produce valid but non-canonical labels (e.g., ‘road’ instead of ‘runway’, or ‘screen’ instead of ‘arcade machine’), resulting in unfair penalization.
@@ 4.2 Open-Vocabulary Segmentation Performance Analysis
MSeg (B) [26] 70 (M) 33.7 32.6 19.1 * 73.4 43.4 * 29.6 33.4 * *
GroupViT (S) [64] 44 (M) - - * * 52.3 22.4 * * * - *
LSeg+ (B) [28] 112 (M) - - 18.0 3.8 * 46.5 7.8 * * - *
ZegFormer (B) [15] 60 (M) - - * 8.1 80.7 * * * * - -
OpenSeg (B) [19] 86 (M) - - 26.4 8.1 70.2 44.8 11.5 * * - *
OVSeg (B) [34] 0.6 (M) - - 29.6 9.0 94.5 55.7 12.4 * * - *
MaskCLIP (L) [16] 428 (M) 15.1 6.0 23.7 8.2 * 45.9 10.0 * * * *
OpenSeeD (L) [73] 39 (M) 19.7 15.0 23.4 * * * * * * * *
X-Decoder-Seg+ (B) [80] 28 (M) 16.9 9.5 23.8 4.6 97.8 64.7 12.1 32.2 35.1 33.8 18.5
X-Decoder (L) [80] 38 (M) 21.8 13.1 29.6 9.2 97.7 64.0 16.1 43.0 49.5 39.5 29.7
APE-B (L) [52] 42 (M) 26.4 23.5 29.0 9.2 95.8 58.3 21.0 * * * *
ESC-Net [27] 451 (M) - - 41.8 18.1 98.3 65.6 27.0 * * - *
@@
Table 2: Oracle-Prompts evaluation of open-vocabulary segmentation models. We report the state-of-the-art (SOTA) model X-Decoder [80]’s performance under both evaluation protocols. Other methods are omitted either because: 1) they are not SOTA, or 2) they do not support oracle-prompts evaluation.
@@
X-Decoder (L) [80] Global-Matching (default) 29.6 9.2 97.7 64.0 16.1 43.0 29.7
@@
Zero-shot open-vocabulary transfer. OpenWorldSAM generalizes out-of-the-box to a broad set of segmentation tasks without any weight adaptation. As shown in Table 1, it achieves state-of-the-art performance across almost all datasets and evaluation metrics. Its performance consistently surpasses strong baselines such as X-Decoder and APE, despite using only 4.5M trainable parameters. On ADE20K-857, OpenWorldSAM achieves 33.1% mIoU, outperforming the previous best (X-Decoder) by +23.9 absolute points (9.2 → 33.1). On PASCAL Context-459, it achieves 47.5% mIoU, improving over APE’s 21.8% by +25.7 points, and on ScanNet-40, it reaches 55.6% mIoU, a +25.9 point improvement over X-Decoder’s 29.7%. On AP score we under-perform APE, which included extra detection datasets, e.g., Objects365 [51], in their training recipe for better localization.
@@
Oracle-Prompts evaluation. As SAM-style models are designed for interactive segmentation, oracle prompts closely reflect practical use cases—such as human-in-the-loop annotation, robotic object search, or dynamic UI feedback. To fairly compare with the state-of-the-art generalist model X-Decoder [80], we also evaluate it under oracle prompts: we restrict its vocabulary to the ground-truth classes for each image. As shown in Table 2, OpenWorldSAM continues to outperform even under these controlled conditions. Notably, on large-vocabulary datasets such as ADE20K-857 and PASCAL Context-459, OpenWorldSAM achieves 33.1% and 47.5% mIoU, surpassing X-Decoder by +4.0 and +5.2 points, respectively. This highlights our model’s superior language grounding ability in long-tailed, fine-grained category distributions. On smaller datasets like PASCAL Context-59 and PASCAL VOC-20, where most categories overlap with COCO, X-Decoder slightly outperforms our model (75.5% vs. 73.7% mIoU and 98.1% vs. 98.0%), suggesting it benefits more from class memorization in such settings. Moreover, Figure 5 illustrates that global matching often fails despite producing correct masks. Conditioning on oracle prompts significantly reduces this ambiguity, highlighting the robustness of our evaluation protocol and the effectiveness of prompt-based segmentation.